OpenStudy (anonymous):

If ||u||= 4 , ||v||=3 and u*v = 3, find ||u+2v||

OpenStudy (anonymous):

expand ||u + 2u||

OpenStudy (kinggeorge):

For this, you'll need to use a particular property of the norm of a vector. Specifically, if \(\vec{w}\) is a vector, then \[||\vec{w}||=\sqrt{w\cdot w}\]Replace \(\vec{w}\) with \(u+2v\), and suddenly everything will fall out.

OpenStudy (kinggeorge):

One more thing you might want to remember, is that if you take \((u+2v)\cdot(u+2v)\), you can expand that using FOIL.

OpenStudy (anonymous):

I worked it out so that I had \[u*u + (2v*u)(2v*u)\] but I feel like that's going in the completely wrong direction. Or is that what I want to end up getting? I worked it through and got \[\sqrt{96}\] as an answer

OpenStudy (kinggeorge):

That doesn't look right. Let's go through this step by step.\[\begin{aligned} ||u+2v||&=\sqrt{(u+2v)\cdot(u+2v)}\\ &=\sqrt{u\cdot u+4u\cdot v+4v\cdot v}\\ &=\sqrt{||u||^2+4(u\cdot v)+4||v||^2}. \end{aligned}\]Now that we've reached this point, we just have to plug in numbers.\[\begin{aligned} ||u+2v||&=\sqrt{16+4\cdot3+9}\\ &=\sqrt{16+12+9}\\ &=\sqrt{37}. \end{aligned}\]Did you follow all of these steps?

OpenStudy (anonymous):

Yes, I see it now! Thanks so much!